Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

W e     h a v e     P ( A ) = 2 5 ,     P ( B ) = 1 3     a n d     P ( C ) = 1 2 P ( A ∩ C ) = 1 5     a n d     P ( B ∩ C ) = 1 4 ∴     P ( C / B ) = P ( B ∩ C ) P ( B ) = 1 / 4 1 / 3 = 3 4 P ( A ' ∩ C ' ) = 1 − P ( A ∪ C ) = 1 − [ P ( A ) + P ( C ) − P ( A ∩ C ) ]                                                     = 1 − [ 2 5 + 1 2 − 1 5 ] = 1 − 7 1 0 = 3 1 0 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t i e s     a r e     3 4     a n d     3 1 0 .

Thumbs Up IconUpvote Thumbs Down Icon
New question posted•a year ago
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

W e     h a v e     P ( A ) = 1 2 ,     P ( B ) = 1 3     a n d     P ( A ∩ B ) = 1 4                                       P ( A ' ) = 1 − 1 2 = 1 2 ,     P ( B ' ) = 1 − 1 3 = 2 3 P ( A ' ∩ B ' ) = 1 − P ( A ∪ B ) = 1 − [ P ( A ) + P ( B ) − P ( A ∩ B ) ]                                                     = 1 − [ 1 2 + 1 3 − 1 4 ] = 1 − [ 6 + 4 − 3 1 2 ] = 1 − 7 1 2 = 5 1 2 ( i )     P ( A / B ) = P ( A ∩ B ) P ( B ) = 1 / 4 1 / 3 = 3 4 ( i i )     P ( B / A ) = P ( A ∩ B ) P ( A ) = 1 / 4 1 / 2 = 1 2 ( i i i )     P ( A ' / B ) = P ( A ' ∩ B ) P ( B ) = P ( B ) − P ( A ∩ B ) P ( B ) = 1 − P ( A ∩ B ) P ( B )                                                               = 1 − 1 / 4 1 / 3 = 1 − 3 4 = 1 4 ( i v )     P ( A ' / B ' ) = P ( A ' ∩ B ' ) P ( B ' ) = 5 / 1 2 2 / 3 = 5 1 2 × 3 2 = 5 8

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

We  know  that  random  variable  X  takes  values  0, 1, 2, 3,…, n  is  said  to  be  binomial  distribution h a v i n g     p a r a m e t e r s     n     a n d     p ,     i f     t h e     p r o b a b i l i t y     i s     g i v e n     b y P ( X = r ) = C r n p r q n − r ,     w h e r e     q = 1 − p     a n d     r = 0 ,   1 ,   2 ,   3 , … S i m i l a r l y ,     i n     c a s e     o f     t o s s i n g     a     c o i n     3     t i m e s , n = 3     a n d     X     h a s     t h e     v a l u e s     0 ,   1 ,   2 ,   3 , …     w i t h     p = 1 2 ,     q = 1 2 . H e n c e ,     i t     i s     s a i d     t o     h a v e     a     b i n o m i a l     d i s t r i b u t i o n .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

T w o     d i c e     a r e     t h r o w n     t o g e t h e r ∴ n ( S ) = 3 6                     E = A     t o t a l     o f     4 = { ( 2 , 2 ) , ( 1 , 3 ) , ( 3 , 1 ) }           ∴ n ( E ) = 3                     F = A     t o t a l     o f     9     o r     m o r e                               = { ( 3 , 6 ) , ( 6 , 3 ) , ( 5 , 4 ) , ( 4 , 5 ) , ( 5 , 5 ) , ( 4 , 6 ) , ( 6 , 4 ) , ( 5 , 6 ) , ( 6 , 5 ) , ( 6 , 6 ) }           ∴ n ( F ) = 1 0                   G = A     t o t a l     d i v i s i b l e     b y     5                               = { ( 1 , 4 ) , ( 4 , 1 ) , ( 2 , 3 ) , ( 3 , 2 ) , ( 4 , 6 ) , ( 6 , 4 ) , ( 5 , 5 ) }           ∴ n ( G ) = 7 H e r e ,     w e     s e e     t h a t     ( E ∩ F ) = ?     a n d     ( E ∩ G ) = ? a n d             ( F ∩ G ) = { ( 4 , 6 ) , ( 6 , 4 ) , ( 5 , 5 ) } ∴                   n ( F ∩ G ) = 3     a n d     ( E ∩ F ∩ G ) = ? ∴                                     P ( E ) = n ( E ) n ( S ) = 3 3 6 = 1 1 2                                           P ( F ) = n ( F ) n ( S ) = 1 0 3 6 = 5 1 8     a n d     P ( G ) = n ( G ) n ( S ) = 7 3 6                           P ( F ∩ G ) = 3 3 6 = 1 1 2     a n d     P ( F ) . P ( G ) = 5 1 8 . 7 3 6 = 3 5 6 4 8 Since,  P(F∩G)≠P(F).P(G) H e n c e ,     t h e r e     i s     n o     p a i r     o f     i n d e p e n d e n t     e v e n t s .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

L e t     r e d     m a r b l e s     b e     r e p r e s e n t e d     w i t h     R     a n d     b l a c k     m a r b l e s     b e     r e p r e s e n t e d     w i t h     B . T h e     f o l l o w i n g     t h r e e     c o n d i t i o n s     a r e     p o s s i b l e ,     i f     a t l e a s t     o n e     o f     t h e     t h r e e     m a r b l e s d r a w n     b e     b l a c k     a n d     f i r s t     m a r b l e     i s     r e d . ( i )         E 1 = I I     b a l l     i s     b l a c k     a n d     I I I     i s     r e d ( i i )     E 2 = I I     b a l l     i s     b l a c k     a n d     I I I     i s     a l s o     b l a c k ( i i i ) E 3 = I I     b a l l     i s     r e d     a n d     I I I     i s     b l a c k ∴               P ( E 1 ) = P ( R 1 ) . P ( B 1 / R 1 ) . P ( R 2 / R 1 B 1 ) = 5 8 . 3 7 . 4 6 = 6 0 3 3 6 = 5 2 8                     P ( E 2 ) = P ( R 1 ) . P ( B 1 / R 1 ) . P ( B 2 / R 1 B 1 ) = 5 8 . 3 7 . 2 6 = 3 0 3 3 6 = 5 5 6 a n d     P ( E 3 ) = P ( R 1 ) . P ( R 2 / R 1 ) . P ( B 1 / R 1 R 2 ) = 5 8 . 4 7 . 3 6 = 6 0 3 3 6 = 5 2 8 ∴                   P ( E ) = P ( E 1 ) + P ( E 2 ) + P ( E 3 ) = 5 2 8 + 5 5 6 + 5 2 8 = 2 5 5 6 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y     i s     2 5 5 6 .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

W e     k n o w     t h a t : A ∪ B     d e n o t e s     t h a t     a t l e a s t     o n e     o f     t h e     e v e n t s     o c c u r s     a n d A∩B  denotes  that  the  two  events  occurs  simultaneously. S o ,               P ( A ∪ B ) = P ( A ) + P ( B ) − P ( A ∩ B ) ⇒                                                 0 . 6 = P ( A ) + P ( B ) − 0 . 3 ⇒                                                 0 . 9 = P ( A ) + P ( B ) ⇒                                                 0 . 9 = 1 − P ( A ¯ ) + 1 − P ( B ¯ ) ⇒ P ( A ¯ ) + P ( B ¯ ) = 2 − 0 . 9 = 1 . 1 H e n c e ,     t h e     r e q u i r e d     a n s w e r     i s     1 . 1

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

A c c o r d i n g     t o     t h e     s o l u t i o n     o f     Q . 1 ,     w e     h a v e A = { ( 1 , 1 ) , ( 2 , 2 ) , ( 3 , 3 ) , ( 4 , 4 ) , ( 5 , 5 ) , ( 6 , 6 ) } ∴ n ( A ) = 6     a n d     n ( S ) = 6 × 6 = 3 6 S o ,         P ( A ) = n ( A ) n ( S ) = 6 3 6 = 1 6 a n d                       B = { ( 4 , 6 ) , ( 6 , 4 ) , ( 5 , 5 ) , ( 5 , 6 ) , ( 6 , 5 ) , ( 6 , 6 ) }                         n ( B ) = 6     a n d     n ( S ) = 6 × 6 = 3 6 ∴                   P ( B ) = n ( B ) n ( S ) = 6 3 6 = 1 6                         A ∩ B = { ( 5 , 5 ) , ( 6 , 6 ) } ∴ P ( A ∩ B ) = 2 3 6 = 1 1 8 T h e r e f o r e ,     i f     A     a n d     B     a r e     i n d e p e n d e n t ,     t h e n           P ( A ∩ B ) = P ( A ) . P ( B ) ⇒                                   1 1 8 ≠ 1 6 × 1 6 = 1 3 6 ⇒                                   1 1 8 ≠ 1 3 6 H e n c e ,     A     a n d     B     a r e     n o t     i n d e p e n d e n t     e v e n t s .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Here,  Sample  space  S={(1,2),(2,1),(1,3),(3,1),(2,3),(3,2),(1,4),(4,1),(1,5),(5,1),(2,4)(4,2),(2,5),(5,2),(3,4),(4,3),(3,5),(5,3),(5,4),(4,5)}∴  n(S)=20Let  X  be  the  random  variable  denoting  the  sum  of  the  numbers  on  two  cards  drawn.∴               X=3,4,5,6,7,8,9So,  P(X=3)=220                          P(X=4)=220        P(X=5)=420                           P(X=6)=420        P(X=7)=420                            P(X=8)=220         P(X=9)=220∴      the  mean  E(X)=∑i=1nXiPi               =3×220+4×220+5×420+6×420+7×420+8×220+9×220                =620+820+2020+2420+2820+1620+1820=12020=6∴                        E(X2)=∑i=1nPiXi2                =9×220+16×220+25×420+36×420+49×420+64×220+81×220                =1820+3220+10020+14420+19620+12820+16220=78020=39∴              Variance(X)=E(X2)−[E(X)]2                                              =39−(6)2=39−36=3

Thumbs Up IconUpvote Thumbs Down Icon